i need help to figure this out.
Simplify
m(2+n-m)+3(3n+m^2-1)

Answer by rothauserc(4718)

m(2+n-m)+3(3n+m^2-1) =
2m + mn -m^2 + 9n +3m^2 -3 =
2m^2 +mn +2m +9n -3 =
2m(m +1) +n(m +9) -3